OpenStudy (anonymous):

For the following equation, log(6.9-y) - log(y) = log(1.1+x4) - 5log(x) Express y explicitly as a function of x. Hence, determine the value of y when x = 3.4, giving your answer to 3 decimal places.

OpenStudy (campbell_st):

so using log laws \[\log(6.9-y)-\log(y) = \log(1.1 + x^4) - 5\log(x)\] \[\log{((6.9-y)/y)} = \log((1.1 + x^4)/x^5) \] \[\log(6.9/y - y/y) = \log(1.1/x^5 + x^4/x^5)\] find the exponent of each side \[6.9/y - 1 = 1.1x^{-5} + x^{-1}\] take the reciprocal of both sides \[y/6.9 - 1 = x^5/1.1 - x\] \[y = 6.9(x^5/1.1 + x + 1)\]

OpenStudy (campbell_st):

I'll let you do the substitution of 3.4 to find y

OpenStudy (anonymous):

thanks for your help!

jhonyy9 (jhonyy9):

log(6,9-y)/y =log(1,1 +x^4)/x^5 so than (6,9-y)/y =(1,1 +x^4)/x^5 x^5(6,9-y)=y(1,1 +x^4) 6,9x^5 -yx^5 =y(1,1 +x^4) 6,9x^5 =y(1,1 +x^4)+yx^5 6,9x^5 =y(1,1 +x^4 +x^5) y=(6,9x^5)/(1,1 +x^4 +x^5) when x=3,4 so than y=(6,9(3,4)^5)/(1,1 +(3,4)^4 +(3,4)^5) =(6,9(454,354))/(1,1 +133,633 +454,354)= =3135,044/589,087 =5,321 3,4^4 = 133,633 3,4^5 =454,354 so the answer wann to be 5,321

OpenStudy (anonymous):

cambell_st, i think your answer is wrong.. i tried and manage to get the same answer as jhonyy